OpenStudy (anonymous):

In Oregon, 3/5 of bottles are recycled each year. A bottling company produces 1.2 million bottles and figures that x years later the number U of those sill being used is given by U(x)=1200270(3/5)^x Of those 1.2 million bottles, how many are still in use 8 years later?

OpenStudy (anonymous):

i am not sure is the andser is 20160 or 2520. I have set U(x)=1200270(3/5)^x to look like U(8)=1200270(3/5)^8 and calculated for 2520. However, i am not sure if you have to divide by 8 on both sides because if I didnt i would get 20160. This is a multiple choice problem and the option are A. 20160 B. 9873 C. 1126024 and D. None of the Above so i am not sure if it is A or D..... PLEASE HELP

OpenStudy (anonymous):

@RadEn @Mertsj

OpenStudy (mertsj):

\[U(x)=1200270(\frac{3}{5})^8=1200270(.01679616)=20160\]